# Self-Calibrating Gaussian Splatting for Large Field of View Reconstruction

### v1

@@ 7 Additional Ablation Studies
(a) Lego | (b) Car | ||
(c) Rover | (d) Spaceship |
8 Additional Results
FisheyeNeRF Scenes [34].
We provide more rendering results in this section. We render both fisheye and perspective view on FisheyeNeRF dataset [34]. In Fig. 18, we render fixed the view direction and camera location of fisheye rendering and extend the FOV of perspective rendering. As for scenes such as cube, chairs, and flowers, we can see the straight lines are nicely recoverred during the reconstruction. The lines on the wall back of cube and lines on the windows are good proof that our self-calibration system accurately model lens distortion.
Synthetic Radial Objects.
We apply moderate radial distortion to our synthetic dataset and reconstruct several object scenes. After training, we can render undistorted images. The perspective rendering increases the FOV while maintaining the same camera extrinsics. As shown in Fig. 19, distorted edge lines, such as those on the Lego and Car objects, are correctly recovered into straight lines, demonstrating the capability of our hybrid field to model radial distortion effectively.
@@ 9 Implementation Details
Our implementation is based on the codebase by Gaussian Splatting [37] and gsplat [64]. The invertible ResNet is constructed using FrEIA [3]. We follow Kerbl et al. [37] to select hyper-parameters for optimizing 3D Gaussians. We also adopt the implementation of MCMC densification [38]. We notice that compared with vanilla densification, MCMC can help to remove floaters by using opacity thresholding to relocate dead Gaussians. While the final quantitative number in the test set remains roughly unchanged, applying the MCMC technique produces fewer visual floaters in novel viewpoints. For high-resolution scene captures, such as Backyard and Office, we also use bilateral grid and anti-anlising [66] for better quality. As explained in Fig. 16, optimizing our hybrid field is essential for successful self-calibration. We use Adam [39] for the invertible ResNet. The initial learning rate for the invertible ResNet is set at 1e-5 and decreases to 1e-7 for FisheyeNeRF [34]. The final learning rate for real-world captures is 1e-8, including Studio, Garden, Backyard in Figure 14, and more complex real-world captures like Office Fig. 13. The learning rate for Mitsuba indoor synthetic scenes is set to 1e-8, while objects are 1e-7. After estimating distortion parameters from COLMAP, we uniformly sample points following the estimated distortion parameters to initialize the invertible ResNet, which typically takes approximately 1 minute to complete. All experiments are conducted on a single NVIDIA GeForce RTX 3090.
@@ 10 Failure Cases and Limitations
Real-world outdoor captures often include the sky. Reconstructing the sky poses challenges due to moving clouds and the large uniform region of blue and white without textures. The 3DGS [37] method tends to assign large Gaussians to the sky, resulting in artifacts when rendering novel views. Occasionally, some large Gaussians leak into the scene’s center, appearing as a thin film in front of the camera. Similarly, for indoor scenes, regions with uniform textures, such as colored walls, present challenges. These textureless walls are often represented by Gaussians with large covariant matrix, causing similar rendering artifacts as observed with the sky.
The Gaussian sorting we propose alleviates the intensity discontinuities at the boundaries of cubemap faces caused by the multiple projections of a single Gaussian. However, since the projection of 3D covariance follows Equation 5 in the main paper, identical 3D Gaussians can still result in different 2D covariates on different faces. This issue can be addressed by implementing smoother transitions between projection faces.
As mentioned in Section 6, we rely on COLMAP [58] to obtain the extrinsic parameters of the camera pose. However, the projection models defined in COLMAP cannot handle the peripheral regions projected from the back of the image plane, leading to complete failure. We circumvent this by manually cropping the raw images into smaller FOVs before inputting them into COLMAP. It would be interesting future direction to improve the reconstruction pipeline to handle this issue automatically.
Finally, we do not account for the entrance pupil shift phenomenon commonly observed in fisheye lenses. This effect is distinct from the lens distortion we are currently modeling. As a result, our method still struggles with such cameras, as shown in Figure 13. While entrance pupil shift is negligible for distant scenes, it can cause splat misalignments in near-field scenes (e.g., the blurry sphere surface in the Office scene shown in the video), as the shift can reach up to half a centimeter for full-frame lenses. It remains an exciting direction to study how to model such lens effects to further improve reconstruction qualities.

### v2

@@ 5.8 Qualitative Results in Undistorted Rendering
We provide additional rendering results in this section. We render both fisheye and perspective views on the FisheyeNeRF dataset [36]. In Fig. 21, we fix the view direction and camera location for fisheye rendering and extend the FOV for perspective rendering. In scenes such as Cube, Chairs, and Flowers, we observe that straight lines are accurately recovered during reconstruction. The lines on the wall behind the Cube and the window frames serve as strong evidence that our self-calibration system precisely models lens distortion.
@@ 6 Implementation Details
Our implementation is based on the codebase from Gaussian Splatting [38] and gsplat [63]. We use the same loss function as 3DGS for training [38]. The invertible ResNet is constructed using FrEIA [3]. We follow Kerbl et al. [38] to select hyperparameters for optimizing 3D Gaussians. We also adopt the implementation of MCMC densification [39]. Compared with vanilla densification, MCMC helps remove floaters by using opacity thresholding to relocate dead Gaussians. While the final quantitative results on the test set remain largely unchanged, applying the MCMC technique reduces visual floaters in novel viewpoints. For high-resolution scene captures such as Backyard and Office, we also use bilateral grids and anti-aliasing [65] for improved quality.
As explained in Fig. 12, optimizing our hybrid field is essential for successful self-calibration. We use Adam [40] to train the invertible ResNet. The initial learning rate for the invertible ResNet is set to 1e-5 and gradually decreases to 1e-7 for FisheyeNeRF [36]. The final learning rate for real-world captures is 1e-8, including Studio, Garden, and Backyard in Figure 18, as well as more complex real-world captures such as Office Fig. 17. The learning rate for Mitsuba indoor synthetic scenes is set to 1e-8, while for object-centric scenes, it is 1e-7. All experiments are conducted on a single NVIDIA GeForce RTX 3090.
@@ 7 Failure Cases and Limitations
Real-world outdoor captures often include the sky. Reconstructing the sky poses challenges due to moving clouds and the large uniform regions of blue and white without textures. The 3DGS [38] method tends to assign large Gaussians to the sky, resulting in artifacts when rendering novel views. Occasionally, some large Gaussians leak into the scene’s center, appearing as a thin film in front of the camera. Similarly, for indoor scenes, regions with uniform textures, such as colored walls, present challenges. These textureless walls are often represented by Gaussians with large covariance matrices, causing similar rendering artifacts as observed with the sky.
The Gaussian sorting we propose alleviates the intensity discontinuities at the boundaries of cubemap faces caused by the multiple projections of a single Gaussian. However, since the projection of 3D covariance follows the equation in Sec. 3.1 of the main paper, identical 3D Gaussians can still result in different 2D covariances on different faces. This issue can be addressed by implementing smoother transitions between projection faces, such as spherical projection.
Finally, we do not account for the entrance pupil shift phenomenon commonly observed in fisheye lenses. This effect is distinct from the lens distortion we are currently modeling. As a result, our method still struggles with such cameras, as shown in Figure 17. While entrance pupil shift is negligible for distant scenes, it can cause splat misalignments in near-field scenes (e.g., the blurry sphere surface in the Office scene shown in the video), as the shift can reach up to half a centimeter for full-frame lenses. It remains an exciting direction to study how to model such lens effects to further improve reconstruction quality.